OpenStudy (anonymous):

If the sin 60° is √3/2, then the cos ____° = _____.

OpenStudy (anonymous):

@calculusxy

zepdrix (zepdrix):

\(\Large\bf \color{#008353}{\text{Hey, Welcome to OpenStudy! :)}}\)

zepdrix (zepdrix):

I'm trying to remember the right word to explain this one. The cosine of the an angle is equal to the `compliment` of the sine of that angle. So like:\[\Large\bf\sf \sin(90^o)\quad=\quad \cos(90^o-90^o)\quad=\quad \cos(0^o)\]\[\Large\bf\sf \sin(60^o)\quad=\quad \cos(90^o-60^o)\quad=\quad \cos(30^o)\]

zepdrix (zepdrix):

\[\Large\bf\sf \sin(60^o)\quad=\quad \cos(30^o)\]If,\[\Large\bf\sf \sin(60^o)\quad=\quad \frac{\sqrt3}{2}\]then,\[\Large\bf\sf \cos(30^o)\quad=\quad ?\]

OpenStudy (anonymous):

wasnt that the original question?

OpenStudy (anonymous):

i know cos(30˚) but the square root threw me off

zepdrix (zepdrix):

They just want you to recognize that sin 60° = cos 30˚ So you should get the same answer for cos 30˚ that sin 60˚ gave you.

OpenStudy (anonymous):

ohh that makes so much more sense.. thx..